### v1

# A. I. Bolotnikov, A. A. Irmatov

Abstract
This article presents examples of an application of the finite field method for the computation of the characteristic polynomial of the matching arrangement of a graph. Weight functions on edges of a graph with weights from a finite field are divided into proper and improper functions in connection with proper colorings of vertices of the matching polytope of a graph. An improper weight function problem is introduced, a proof of its NP-completeness is presented, and a knapsack-like public key cryptosystem is constructed based on the improper weight function problem.
Introduction
In [1] the matching arrangement of hyperplanes $MA(G)$ for a graph $G(V,E),\\ |E|=n$ without loops and parallel edges was introduced as the arrangement that contains a hyperplane with the equation $x_{i_{1}}-x_{i_{2}}+...+(-1)^{k+1}x_{i_{k}}=0$ for every sequence of edges $(e_{i_{1}},...,e_{i_{k}})$ that form either a simple path of arbitrary length or a simple cycle of even length in $G$. The matching arrangement has the following property: if you arbitrarily select two vectors from the same region of the arrangement, use each vector in turn as a weight vector on the edges of graph $G$, and find the maximum-weight matching for each vector, then for both vectors, the maximum matching will be the same [1]. In addition, there is a one-to-one correspondence between the regions of the matching arrangement and the LP-orientations of the matching polytope, a special polytope used in solving the maximum matching problem using linear programming methods [2].
In [3], Thomas Zaslavsky showed that the number of regions in a hyperplane arrangement depends on the partially ordered set of intersections of the hyperplanes in the arrangement. Zaslavsky expressed this number through the characteristic polynomial of the hyperplane arrangement. Thus, if we calculate the characteristic polynomial of a hyperplane arrangement, we can find its number of regions.
The results on the properties of the characteristic polynomial of a matching arrangement, as well as on calculating the characteristic polynomial for the case when graph $G$ is a tree, are also presented in [1]. In this paper, we will consider examples of calculating the characteristic polynomial of a matching arrangement using the finite field method. This method was developed by Athansiadis [4] based on the results of Crapo and Rota [5], and allows us to calculate the characteristic polynomial of a hyperplane arrangement if all equations of the hyperplanes in the arrangement have rational coefficients. The method is based on the following result [4, Theorem 2.1]:
Let A be any subspace arrangement in $R^{n}$ defined over the integers and q be a large enough prime number. Then $\chi_{A}(q)=|F_{q}^{n}\setminus\cup A|$ Equivalently, identifying $F_{q}^{n}$ with $\{0,1,...,q-1\}^{n}$, $\chi_{A}(q)$ is the number of points with integer coordinates in the cube $\{0,1,...,q-1\}^{n}$ which do not satisfy mod q the defining equations of any of the subspaces in A.
The second section of this paper presents results on calculating the characteristic polynomials of the matching arrangement for some families of graphs using the finite field method.The definitions of a hyperplane arrangement, an intersection lattice and a characteristic polynomial that were used in the first part can be found in [6], which also contains the proof of the deletion-restriction lemma and the description of the finite field method that uses powers of prime.
In the third section, we introduce the definition of a proper and improper weight function, state the improper weight function problem, prove its NP-completeness, and describe the relationship between proper weight functions and the matching arrangement. The NP-completeness is proved with the use of 3-satisfiability, or 3-SAT problem. All the necessary definitions regarding the 3-SAT problem, as well as proof of its NP-completeness, can be found in [7].
The forth section of the paper provides a theoretical description of the Bolotnikov-Irmatov cryptosysyem that uses the NP-completeness of the improper weight function problem.
The results from the second and the third sections are obtained by A. I. Bolotnikov. The cryptosystem in the forth section is developed by A. I. Bolotnikov and A. A. Irmatov.
The characteristic polynomial of the matching arrangement for some classes of graphs.
Proposition 1.
Let $G$ be a graph that consists of one cycle of length $n$, and let $\chi_{MA(G)}(x)$ be the characteristic polynomial of the matching arrangement $MA(G)$. Then
  • if $n$ is even, then $\chi_{MA(G)}(x)=(x-1)(x-n)^{n-1}$.
  • if $n$ is odd, then $\chi_{MA(G)(x)}=(x-1)(x-3)(x-5)...(x-(2n-3))(x-(n-1))$.
Proof.
Let graph G consist of one cycle of length n. We arbitrarily choose vertex A in graph G and number the edges clockwise, counting the edges from vertex A.
Figure 1: A cycle of length n
Let us first consider the case when n is even. Consider the linear operator $W:R^{n}\rightarrow R^{n}$ with the following matrix. The main diagonal of the matrix, as well as the next diagonal to the right, alternate between 1 and -1 (starting with 1), and there are 0s in all other places.
$\left(\begin{matrix}1&1&0&0&0&0\\ 0&-1&-1&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&-1&-1&0\\ 0&0&0&0&1&1\\ 0&0&0&0&0&-1\\ \end{matrix}\right)$
The vector $t=(1,-1,...,(-1)^{k+1},0,...,0)$ is the normal vector to the hyperplane from the arrangement $MA(G)$, which corresponds to a path of length k leaving A and going clockwise. The operator W maps this vector to vector $d_{k}=(0,...,0,1,0,...,0)$, in which 1 is at the $k$-th position. For an arbitrary vertex B of graph G, the vector $t_{B}=(1,-1,...,(-1)^{r+1},0,...,0)$, where r is the length of the path from A to B clockwise, is normal to the hyperplane corresponding to the path from A to B in a clockwise direction. $t_{A}=(1,-1,...,1,-1)$ will denote the normal to the hyperplane corresponding to the entire cycle. Consider an arbitrary path in graph G that leads clockwise from vertex B to vertex C. If $B=A$, then the vector $t_{C}$ will be normal to the hyperplane corresponding to this path. If $C=A$ or the path does not pass through A, then $t_{C}-t_{B}$ will be normal to the corresponding hyperplane. If the path passes through A but does not start or end there, then $t_{A}-t_{B}+t_{C}$ will be normal to the corresponding hyperplane. Operator W maps $t_{B}$ to $d_{r}$, $t_{C}-t_{B}$ to $d_{k_{2}}-d_{k_{1}}$, where $k_{2}>k_{1}$, and $t_{A}-t_{B}+t_{C}$ to $d_{n}-d_{k_{1}}+d_{k_{2}}$, where $k_{1}>k_{2}$.
Consider the hyperplane arrangement S that consists of the following hyperplanes:
  1. 1.
    $\forall 1\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n:x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}+x_{i_{1}}=0$
Note that since there is an invertible linear operator W that maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of S , their intersection lattices are isomorphic and their characteristic polynomials are equal. We will look for the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
  1. 1.
    $\forall 1\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n:x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}+x_{i_{1}}=0$.
We will find the number $\chi_{S}(p)$ combinatorially. First, we choose a value for $x_{n}$. This can be done in $p-1$ ways. Let’s say we have chosen $x_{n}=k$, $k\neq 0$. Then the remaining $x_{i}$ should be selected so that they do not satisfy any of the following conditions:
  1. 1.
    $\forall 1\leq i\leq n-1:x_{i}=0$
  2. 2.
    $\forall 1\leq i\leq n-1:x_{i}=k$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}-x_{i_{1}}=0$
  4. 4.
    $\forall 1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}-x_{i_{1}}=k$.
Let’s choose a weak ordered partition $\pi=(B_{1},B_{2},....,B_{p-n})$ of the set $\{1,2,3,..,n-1\}$ into $p-n$ blocks (“weak” means that some of the blocks can be empty), that is $\bigcup B_{i}=\{1,2,...,n\}$ and $B_{i}\cap B_{j}=\emptyset,i\neq j$. Arrange all the elements of $F_{p}$ in a row as follows: $0,k,2k,...,(p-1)k$. The first two elements of this row are “locked”: none of the $x_{i}$ can take these two values. Starting with the third element in this row, elements of $B_{1}$ are placed one by one in the descending order onto the elements of the row without any gaps. Once all elements of $B_{1}$ are placed, a gap with the size of one element of the row is made, and elements of $B_{2}$ are placed on the folowing elements of the row in the same descending order. We keep placing elements of $B_{i}$ onto the elements of the row in the descending order and making one-element gaps when moving from $B_{i}$ to $B_{i+1}$, until all elements from $\{1,2,3,...,n-1\}$ are placed. Now for every i, $10\leq i\leq n-$ 1, we choose the element of the row, in which we placed i, as $x_{i}$. Conditions $x_{i}=0$ and $x_{i}=k$ aren’t satisfied for any i, because the first two elements of the row were left “locked”. Condition $x_{i_{2}}-x_{i_{1}}=0$ is not satisfied for any $i_{1}$ and $i_{2}$, because we put no more then one element of $\{1,2,3,4,..,n-1\}$ onto each element of the row. Condition $0\leq i_{1}\leq i_{2}\leq_{n}-1:x_{i_{2}}-x_{i_{1}}=k$ is not satisfied for any $i_{1}\leq i_{2}$, because for any $B_{i}$ the elements of $B_{i}$ were placed in the descending order, and there was a one-element gap between elements of $B_{i}$ and $B_{i+1}$. Therefore, a vector $(x_{1},x_{2},...,x_{n-1})$ is constructed for a given partition $\pi$, and it does not satisfy any of the four conditions. Conversely, given a vector $(x_{1},x_{2},...,x_{n-1})$ that does not satisfy any of the four conditions, a placement of $\{1,2,3,...,n-1\}$ onto the row $0,k,2k,...,(p-1)k$ can be recovered, and in turn a partition $\pi$ can be recovered from this placement. This means that the number of vectors that do not satisfy any of the four conditions is equal to the number of weak ordered partitions of $\{1,2,3,...,n-1\}$ into $p-n$ blocks. There are exactly $(p-n)^{n-1}$ of such partitions. And considering the choice of $x_{n}$ on the first step, the resulting formula for the characteristic polynomial of S is $\chi_{S}(p)=(p-1)(p-n)^{n-1}$. Thus, in case when graph G consists of one cycle of even length, $\chi_{MA(G)}(x)=(x-1)(x-n)^{n-1}$.
Now let n be odd. Let $L$ be the following arrangement of hyperplanes: $L=MA(G)\cup\{H\}$, where H is a hyperplane with the equation $x_{1}-x_{2}+x_{3}...-x_{n-1}+x_{n}=0$.
Consider the linear operator $W:R^{n}\rightarrow R^{n}$ with the following matrix. The main diagonal of the matrix, as well as the next diagonal to the right, alternate between 1 and -1 (starting with 1), and there are 0s in all other places.
$\left(\begin{matrix}1&1&0&0&0&0&0\\ 0&-1&-1&0&0&0&0\\ 0&0&1&1&0&0&0\\ 0&0&0&-1&-1&0&0\\ 0&0&0&0&1&1&0\\ 0&0&0&0&0&-1&-1\\ 0&0&0&0&0&0&1\\ \end{matrix}\right)$
The vector $t=(1,-1,...,(-1)^{k+1},0,...,0)$ is the normal vector to the hyperplane from the arrangement $L$, which corresponds to a path of length k leaving A and going clockwise. The operator W maps this vector to vector $d_{k}=(0,...,0,1,0,...,0)$, in which 1 is at the $k$-th position. For an arbitrary vertex B of graph G, the vector $t_{B}=(1,-1,...,(-1)^{r+1},0,...,0)$, where r is the length of the path from A to B clockwise, is normal to the hyperplane corresponding to the path from A to B in a clockwise direction. $t_{A}=(1,-1,...,1,-1,1)$ will denote the normal to $H$. Consider an arbitrary path in graph G that leads clockwise from vertex B to vertex C. If $B=A$, then the vector $t_{C}$ will be normal to the hyperplane corresponding to this path. If $C=A$ or the path does not pass through A, then $t_{C}-t_{B}$ will be normal to the corresponding hyperplane. If the path passes through A but does not start or end there, then $t_{A}-t_{B}-t_{C}$ will be normal to the corresponding hyperplane. Operator W maps $t_{B}$ to $d_{r}$, $t_{C}-t_{B}$ to $d_{k_{2}}-d_{k_{1}}$, where $k_{2}>k_{1}$, and $t_{A}-t_{B}-t_{C}$ to $d_{n}-d_{k_{1}}-d_{k_{2}}$, where $k_{1}>k_{2}$.
Consider the hyperplane arrangement S that consists of the following hyperplanes:
  1. 1.
    $\forall 1\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n:x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}-x_{i_{1}}=0$
Note that since there is an invertible linear operator W that maps the set of normal vectors to hyperplanes of $L$ to the set of normal vectors to hyperplanes of S , their intersection lattices are isomorphic and their characteristic polynomials are equal. We will look for the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
  1. 1.
    $\forall 1\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n:x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}-x_{i_{1}}=0$.
We will find the number $\chi_{S}(p)$ combinatorially. First, we choose a value for $x_{n}$. This can be done in $p-1$ ways. Let’s say we have chosen $x_{n}=k$, $k\neq 0$. Then the remaining $x_{i}$ should be selected so that they do not satisfy any of the following conditions:
  1. 1.
    $\forall 1\leq i\leq n-1:x_{i}=0$
  2. 2.
    $\forall 1\leq i\leq n-1:x_{i}=k$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}-x_{i_{1}}=0$
  4. 4.
    $\forall 1\leq i_{1}\leq i_{2}\leq_{n}-1:x_{i_{2}}+x_{i_{1}}=k$.
Let us consider two possible cases:
Case 1: $\forall 0\leq i\leq n-1x_{i}\neq k/2$. In this case we can consequitively choose values for all $x_{i}$. There are $p-3$ possible values for $x_{1}$. After a value is chosen for $x_{1}$, there are $k-5$ possible values for $x_{2}$: if value t is chosen for $x_{1}$, then we cannot choose t and $k-t$ for $x_{2}$. After a value is chosen for $x_{2}$,there are $p-7$ possible values for $x_{3}$, and so on. In total in this case there are $(p-3)(p-5)(p-7)...(p-(2n-1))$ possible ways to choose values for $x_{1},x_{2},...,x_{n_{1}}$.
Case 2: for some i $x_{i}=k/2$. In this case we start with choosing in n-1 ways, which of $x_{i}$ is equal to $k/2$, and then for the rest of $x_{i}$ the values are chosen the same way, as in case 1.Therefore, in this case there are $(n-1)(p-3)(p-5)...(p-(2n-3))$ ways to choose values for $x_{1},x_{2},...,x_{n_{1}}$.
After adding the numbers from two cases, and considering the choice of the value for $x_{n}$ on the first step, the following formula for the characteristic polynomial of S is obtained: $\chi_{S}(p)=(p-1)(p-3)(p-5)...(p-(2n-3))(p-n)$.
Let $\hat{S}$ be an arrangement $\hat{S}=S\setminus{\hat{H}}$, where $\hat{H}$ is an arrangement with the equation $x_{n}=0$. An invertible linear operator W maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of $\hat{S}$ , which means that their intersection lattices are isomorphic and their characteristic polynomials are equal.
The characteristic polynomial of $\hat{S}$ can be found with the use of the deletion-restriction lemma: $\chi_{\hat{S}}(x)=\chi_{S}(x)+\chi_{S_{\hat{H}}}(x)$, where $S_{\hat{H}}$ is an arrangement of hyperplanes in $\hat{H}$ that consists of intersections of $\hat{H}$ with other hyperplanes of S. The following are the equations of hyperplanes of $S_{\hat{H}}$:
  1. 1.
    $\forall 1\leq i\leq n-1:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}+x_{i_{1}}=0$.
We will find $\chi_{S_{\hat{H}}}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations above. Let us choose values for $x_{i}$ consequitively. There are $p-1$ ways to choose a value for $x_{1}$, there are $p-3$ ways to choose a value for $x_{2}$ after choosing a value for $x_{1}$, and so on. Therefore, $\chi_{S_{\hat{H}}}(p)=(p-1)(p-3)(p-5)...(p-(2n-3))$.
After applying the deletion-restriction lemma, the following result is obtained: in case when G consists of one cycle of odd length, $\chi_{MA(G)(x)}=(x-1)(x-3)(x-5)...(x-(2n-3))(x-(n-1))$.
Proposition 2.
Let graph G consist of two subgraphs that intersect in one vertex - one subgraph is a complete graph $K_{3}$ and the other is the “tail” - a path of an arbitrary length. Then $\chi_{MA(G)}(x)=(x-1)(x-3)(x-4)....(x-(n-2))(x-(n-1))(x-(n-1))(x-n)$.
Proof.
Let graph G consist of two subgraphs that intersect in one vertex - one subgraph is a complete graph $K_{3}$ and the other is the “tail” - a path of an arbitrary length. Let us number the edges of G in the following order: first we number the edges of $K_{3}$, then we number the edges of the tail, starting with the edge that is adjacent to $K_{3}$.
Figure 2: Graph G
Let W be a linear operator $W:R^{n}\rightarrow R^{n}$ with the following matrix:
Figure 3: Matrix of the operator W
Submatrix A is this matrix of a size 3 by 3:
$\left(\begin{matrix}1&0&0\\ 0&1&1\\ 0&0&-1\\ \end{matrix}\right)$
The main diagonal of the submatrix B, as well as the next diagonal to the right, alternates between 1 and -1 (starting with 1), and there are 0s in all other places.
$\left(\begin{matrix}1&1&0&0&0&0&0\\ 0&-1&-1&0&0&0&0\\ 0&0&1&1&0&0&0\\ 0&0&0&-1&-1&0&0\\ 0&0&0&0&1&1&0\\ 0&0&0&0&0&-1&-1\\ 0&0&0&0&0&0&1\\ \end{matrix}\right)$
Let $d_{1}=(1,0,...,0)$, $d_{2}=(0,1,0,...,0)$, $d_{3}=(0,1,-1,0,...,0)$,
$d_{i}=(0,0,0,1,-1,...,(-1)^{i},0,...,0),i>3$. Operator W maps $d_{i}$ into a vector $(0,...,0,1,0,...,0)$, where 1 stands on the i-th place.
Let’s consider three possible cases for paths in G.
Case 1: The path is entirely inside the $K_{3}$ subgraph. There are only 6 paths that belong to this case: $\{e_{1}\},\{e_{2}\},\{e_{3}\},\{e_{1},e_{2}\},\{e_{2},e_{3}\},\{e_{3},e_{1}\}$. $d_{1},d_{2},d_{3}-d_{2},d_{2}-d_{1},d_{3},d_{3}-d_{2}+d_{1}$ are normal vectors to the corresponding hyperplanes.
Case 2: The path is entirely inside the tail. If the path contains $e_{4}$, then $d_{i}$ will be a normal vector to the corresponding hyperplane, otherwise $d_{j}-d_{i},j>i$ will be a normal to the corresponding hyperplane.
Case 3: The path goes from $K_{3}$ into the tail. The following are all the possible cases of the intersection of this path and the $K_{3}$ subgraph. If this intersection is $\{e_{1}\}$,then $d_{i}-d_{1}$ is a normal vector to the corresponding hyperplane. If $\{e_{2}\}$ is the intersection, then $d_{i}-d_{2}$ will be the normal vector. If $\{e_{2},e_{3}\}$ is the intersection, then $d_{i}-d_{3}$ will be the normal vector.If $\{e_{3},e_{1}\}$ is the intersection, then $d_{i}-d_{1}+d_{2}-d_{3}$ is the normal vector.
Let S be an arrangement that consists of the following hyperplanes:
  1. 1.
    $\forall 1\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n,(i_{1},i_{2})\neq(1,3):x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 4\leq i\leq n:x_{i}-x_{1}+x_{2}-x_{3}=0$
  4. 4.
    $x_{1}-x_{2}+x_{3}=0$
Note that since there is an invertible linear operator W that maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of S , their intersection lattices are isomorphic and their characteristic polynomials are equal. We will look for the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
  1. 1.
    $\forall 1\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 1\leq i_{1}\leq i_{2}\leq n,(i_{1},i_{2})\neq(1,3):x_{i_{2}}-x_{i_{1}}=0$
  3. 3.
    $\forall 4\leq i\leq n:x_{i}-x_{1}+x_{2}-x_{3}=0$
  4. 4.
    $x_{1}-x_{2}+x_{3}=0$
Assume that values $x_{1}=k_{1},x_{2}=k_{2},x_{3}=k_{3}$ are already chosen. Then the values for the rest of $x_{i}$ should be chosen, so that none of the following conditions are satisfied:
  1. 1.
    $\forall 4\leq i\leq n:x_{i}=0$
  2. 2.
    $\forall 4\leq i\leq n:x_{i}=k_{1}$
  3. 3.
    $\forall 4\leq i\leq n:x_{i}=k_{2}$
  4. 4.
    $\forall 4\leq i\leq n:x_{i}=k_{3}$
  5. 5.
    $\forall 4\leq i\leq n:x_{i}=k_{3}+k_{1}-k_{2}$
  6. 6.
    $\forall 4\leq i_{1}\leq i_{2}\leq n:x_{i_{2}}-x_{i_{1}}=0$
The number of possible ways to choose the rest of $x_{i}$ depends on the number of distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$.
The value for $k_{1}$ can be chosen in $p-1$ ways. Let us consider the following cases:
  • Case 1: $k_{3}=k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1}$, and we can choose $k-2$ in $p-3$ ways. In this case there are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$ therefore there are $(p-1)(p-3)(p-4)...(p-n)$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 2: $k_{2}=2k_{1},k_{3}=3k_{1}$. There are 4 distinct values among $0,k_{1},k_{2},k_{3},\\ k_{3}+k_{1}-k_{2}$ therefore there are $(p-1)(p-4)(p-5)...(p-n)$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 3: $k_{2}=2k_{1},k_{3}\neq 3k_{1},k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq 3k_{1}$, and we can choose $k_{3}$ in $p-4$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$ , therefore there are $(p-1)(p-4)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case .
  • Case 4: $k_{2}=k_{1}/2,k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq k_{2}-k_{1}$, and we can choose $k_{3}$ in $p-4$ ways. There are 5 distint values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$ therefore there are $(p-1)(p-4)(p-5)(p-6)...(p-(n+1))$ to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 5: $k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2,k_{3}=2k_{2}-k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2$, and we can choose $k_{2}$ in $p-4$ ways.There are 4 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore there are $(p-1)(p-4)(p-4)(p-5)...(p-n)$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
  • Case 6: $k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2,k_{3}\neq 2k_{2}-k_{1},k_{3}\neq k_{1}$. The conditions for choosing a value for $k_{2}$ are $k_{2}\neq 0,k_{2}\neq k_{1},k_{2}\neq 2k_{1},k_{2}\neq k_{1}/2$, and we can choose $k_{2}$ in $p-4$ ways. The conditions for choosing a value for $k_{3}$ are $k_{3}\neq 0,k_{3}\neq k_{1},k_{3}\neq k_{2},k_{3}\neq k_{2}-k_{1},k_{3}\neq 2k_{2}-k_{1}$ , and we can choose $k_{3}$ in $p-5$ ways. There are 5 distinct values among $0,k_{1},k_{2},k_{3},k_{3}+k_{1}-k_{2}$, therefore there are $(p-1)(p-4)(p-5)(p-5)(p-6)...(p-(n+1))$ ways to choose values for $(x_{1},x_{2},...,x_{n})$ in this case.
The value of $\chi_{MA(G)}(p)$ is the sum of numbers from these cases. As a result, if $G$ consists of a $K_{3}$ subgraph and a “tail” subgraph that are connected through a shared vertex, the formula for $\chi_{MA(G)}(x)$ is $\chi_{MA(G)}(x)=(x-1)(x-3)(x-4)....(x-(n-2))(x-(n-1))(x-(n-1))(x-n)$. ∎
Improper weight function problem
Let $G(V,E),|E|=n$ be a graph without loops and parallel edges. Let $F_{p}$ be the group of residue classes modulo a prime number. A weight function $W:E\rightarrow F_{p}^{k}$ is called improper, if there is a sequence of edges $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that form a simple path or a simple even cycle in G, such that
$W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{d-1}W(e_{i_{d}})={\bf 0}\in F_{p}^{k}.$ (1)
Otherwise the weight function is called proper. The choice of names is motivated by the following consideration. If a weight function on the edges of graph $G$ is given, then the weight of a matching can be defined as the sum of the weights of its edges. Furthermore, there is a one-to-one correspondence between the matchings of graph G and the vertices of the matching polytope of graph G [8]. Thus, based on the weight function on the edges of graph G, it is possible to construct a weight function on the vertices of the matching polytope of graph G. An important property of the matching polytope is that any two vertices of this polytope are connected with the edge, if an only if the symmetrical difference of the corresponding matchings consists of exactly one simple path of arbitrary length or one simple cycle of even length [9, Theorem 25.3]. If the original weight function on the edges of the graph was proper, then for the resulting weight function on the vertices of the polytope, any two vertices connected by an edge of the polytope have different weights. Conversely, in case of an improper initial weight function on the edges of the graph, at least one pair of vertices with the same weight connected by an edge will be found in the polytope. In this case the path or the cycle that satisfies (1) is the symmetrical difference of two matchings that correspond to vertices of the same weight in the matching polytope that are connected by an edge.
Let us state the improper weight function problem. For an arbitrary graph $G(V,E)$ without loops and parallel edges, an arbitrary prime number p, an arbitrary natural number k and an arbitrary weight function $W:E\rightarrow F_{p}^{k}$ the problem is to establish, if function W is an improper one.
Theorem 1.
The improper weight function problem is NP-complete.
Proof.
We will first prove that the problem belongs to the complexity class NP. By definition, a recognition problem $A$ with possible answers “Yes” or “No”  belongs to the class NP if, when $A$ has an answer “Yes”, there exists a “hint” $c(A)$, the length of which polynomially depends on the size of $A$, such that the task $(c(A),A)$ is solved in polynomial time.
For a improper weight function problem, let us choose a simple path or a simple even cycle $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that satisfies (1), as a “hint”. With such a hint the solution of a improper weight function problem is reduced to computing $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{d-1}W(e_{i_{d}})$. Since there are no more then n edges in this path, the computation time depends polynomially from n and k.
To show that improper weight function problem is an NP-complete one, we now transfrom an arbitrary 3-satisfiability problem into an improper weight function problem. Let $X=\{x_{1},x_{2},...,x_{n}\}$ be a set of boolean variables and let $U=\{u_{1},u_{2},...,u_{m}\}$ be set of clauses making up an arbitrary instance of 3-satisfiability problem. The graph G for the improper weight function problem will be constructed in the following way. For each variable $x_{i}$, there will be a subgraph $Z_{i}$ of the graph G that consists of one cycle of length 4, with the set of edges $\{a_{i},b_{i},c_{i},d_{i}\}$ and the set of vertices $\{A_{i},B_{i},C_{i},D_{i}\}$, such that $a_{i}=(A_{i},B_{i}),b_{i}=(B_{i},C_{i}),c_{i}=(C_{i},D_{i}),d_{i}=(D_{i},A_{i})$.
Figure 4: Subraph $Z_{i}$
For each clause $u_{j}$, there will be a bipartite subgraph $Y_{j}\cong K_{2,7}$, with $\{P_{j},Q_{j}\}$ and
$\{R_{j,1},R_{j,2},R_{j,3},R_{j,4},R_{j_{5}},R_{j,6},R_{j_{7}}\}$ being two parts of $Y_{j}$.
Figure 5: Subraph $Y_{j}$
To construct the graph G, we connect $Z_{i}$ with $Z_{i+1}$ through a shared vertex, starting with $Z_{1}$: $C_{i}=A_{i+1}$. Then we connect $Z_{n}$ with $Y_{1}$ through a shared vertex: $C_{n}=P_{1}$. Then we connect $Y_{j}$ with $Y_{j+1}$ through a shared vertex, starting with $Y_{1}$: $Q_{i}=P_{i+1}$.
Figure 6: An example for $n=5$ and $m=3$
The constructed graph $G(V,E)$ consists of $13n+8m+$ 1 vertices and $4n+14m$ edges. The next step is to construct the weight function $W:E\rightarrow F_{p}^{k}$. The prime number p can be chosen aribtrarily. Let $k=3n+11m-1$. The values of the weight function $W(e)=(w_{1}(e),....,w_{k}(e))$ will be defined in the following way.
Let $\hat{V}=V\setminus\{A_{1},Q_{m}\}$. $|\hat{V}|=3n+8m-1$. Let all the elements of $\hat{V}$ be numbered: $\hat{V}=\{v_{1},...,v_{3n+8m-1}\}$. For $e\in E$ and $11\leq i\leq 3n+8m-$ 1, let $w_{i}(e)=1$, if $e$ is an adjacent edge to $v_{i}$, and let $w_{i}(e)=0$ otherwise.
Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause.
$w_{3n+8m-1+3\cdot(j-1)+1}(e)=1$, if either $t_{\alpha}=x_{\alpha}$ and $e=(A_{\alpha},B_{\alpha})$, or $t_{\alpha}=\neg x_{\alpha}$ and $e=(A_{\alpha},D_{\alpha})$. $w_{3n+8m-1+3\cdot(j-1)+1}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),\\ (Q_{j},R_{j,3}),(Q_{j},R_{j,4})\}$. Otherwise, $w_{3n+8m-1+3\cdot(j-1)+1}(e)=0$.
$w_{3n+8m-1+3\cdot(j-1)+2}(e)=1$, if either $t_{\beta}=x_{\beta}$ and $e=(A_{\beta},B_{\beta})$, or $t_{\beta}=\neg x_{\beta}$ and $e=(A_{\beta},D_{\beta})$. $w_{3n+8m-1+3\cdot(j-1)+2}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),\\ (Q_{j},R_{j,5}),(Q_{j},R_{j,6})\}$. Otherwise, $w_{3n+8m-1+3\cdot(j-1)+2}(e)=0$.
$w_{3n+8m-1+3\cdot(j-1)+3}(e)=1$, if either $t_{\gamma}=x_{\gamma}$ and $e=(A_{\gamma},B_{\gamma})$, or $t_{\gamma}=\neg x_{\gamma}$ and $e=(A_{\gamma},D_{\gamma})$. $w_{3n+8m-1+3\cdot(j-1)+3}(e)=-1$, if $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,3}),\\ (Q_{j},R_{j,5}),(Q_{j},R_{j,7})\}$. Otherwise, $w_{3n+8m-1+3\cdot(j-1)+3}(e)=0$.
Now function W is completely defined. Because of the way that $w_{i}(e),1\leq i\leq 3n+8-1$, were defined, if a path start or ends in any vertex $v_{j}$ from $V\setminus\{A_{1},Q_{m}\}$, there is exactly one edge $e$ on this path that has $w_{j}(e)\neq 0$. Therefore, it is impossible for a simple path $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that satisfies (1) to have any ends, other then $A_{1}$ and $Q_{m}$. Note that the only cycles of even length in $G$ are cycles of length for inside $Z_{i}$ or $Y_{j}$. It is not possible for a cycle of length 4 inside $Z_{i}$ or $Y_{j}$ to satisfy (1) as well, because with these values for $w_{i}(e),i>3n+8-1$, it is nessesary for every path or even cycle $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ that satisfies (1) and contains an edge in one of $Z_{i}$ to also contain an edge in one of $Y_{j}$, and visa versa. In case of $Z_{i}$ the reason for this is that there is a specific index $t$ that corresponds to a pair $(x_{i},u)$, where $u$ is a clause that contains either $x_{i}$ or $\neg x_{i}$, such that there is exactly one edge $e$ in $Z_{i}$ with $w_{t}(e)\neq 0$. In case of $Y_{j}$, each of the edges $e\in\{(Q_{j},R_{j,1}),(Q_{j},R_{j,2}),(Q_{j},R_{j,3}),(Q_{j},R_{j,4}),(Q_{j},R_{j,5}),(Q_{j},R_{j,6}),(Q_{j},R_{j,7})\}$ has its own distinct vector of values for $(w_{3n+8m-1+3\cdot(j-1)+1}(e),w_{3n+8m-1+3\cdot(j-1)+2}(e),\\ w_{3n+8m-1+3\cdot(j-1)+3}(e))$, and since any cycle of length 4 in $Y_{j}$ contains exatly two of those edges, it can’t satisfy (1). Therefore, the only paths or even cycles that can possibly satisfy (1) are paths that connect $A_{1}$ and $Q_{m}$. Note that k, the size of G and the computation of W are all polynomial from m and n.
Let $\{\alpha_{1},\alpha_{2},...,\alpha_{n}\}$ be a set of values assigned to $X$, such that all clauses $u_{1},u_{2},...,u_{n}$ are satisfied. The current goal is to find a path $(e_{i_{1}},e_{i_{2}},...,e_{i_{d}})$ in G that satisfies (1) knowing $\{\alpha_{1},\alpha_{2},...,\alpha_{n}\}$. Let us consider the following path $(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$ of length $2\cdot(m+n)$ from $A_{1}$ to $Q_{m}$. For $1\leq i\leq n$, if $\alpha_{i}=True$, then $e_{2\cdot i-1}=a_{i},e_{2\cdot i}=b_{i}$, and if $\alpha_{i}=False$, then $e_{2\cdot i-1}=d_{i},e_{2\cdot i}=c_{i}$.
Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause. Let’s consider all possible cases, after values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ are assigned.
  • If $t_{\alpha}=True,t_{\beta}=True,t_{\gamma}=True$, then $e_{2n+2j-1}=(P_{j},R_{j,1}),e_{2n+2j}=(Q_{j},R_{j,1})$.
  • If $t_{\alpha}=True,t_{\beta}=True,t_{\gamma}=False$, then $e_{2n+2j-1}=(P_{j},R_{j,2}),e_{2n+2j}=(Q_{j},R_{j,2})$.
  • If $t_{\alpha}=True,t_{\beta}=False,t_{\gamma}=True$, then $e_{2n+2j-1}=(P_{j},R_{j,3}),e_{2n+2j}=(Q_{j},R_{j,3})$.
  • If $t_{\alpha}=True,t_{\beta}=False,t_{\gamma}=False$, then $e_{2n+2j-1}=(P_{j},R_{j,4}),e_{2n+2j}=(Q_{j},R_{j,4})$.
  • If $t_{\alpha}=False,t_{\beta}=True,t_{\gamma}=True$, then $e_{2n+2j-1}=(P_{j},R_{j,5}),e_{2n+2j}=(Q_{j},R_{j,5})$.
  • If $t_{\alpha}=False,t_{\beta}=True,t_{\gamma}=False$, then $e_{2n+2j-1}=(P_{j},R_{j,6}),e_{2n+2j}=(Q_{j},R_{j,6})$.
  • If $t_{\alpha}=False,t_{\beta}=False,t_{\gamma}=True$, then $e_{2n+2j-1}=(P_{j},R_{j,7}),e_{2n+2j}=(Q_{j},R_{j,7})$.
Now the path $P=(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$ is completely defined. Let $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=(z_{1},z_{2},...,z_{k})\in F_{p}^{k}$. For $11\leq i\leq 3n+8m-$ 1, $z_{i}=0$, because for every vertex $v\in V\setminus\{A_{1},Q_{m}\}$ path P either passes through $v$ without stopping there or doesn’t pass through $v$ at all. For $i>3n+8m-1$, $z_{i}$ corresponds to a pair $(x_{q},u_{p})$, where $u_{p}$ contains either $x_{q}$ or $\neg x_{q}$. There are four possible cases:
  • Case 1: $\alpha_{q}=True$, $u_{p}$ contains $x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=1,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=-1,w_{i}(e)=0$ for every other edge e on P.
  • Case 2: $\alpha_{q}=False$, $u_{p}$ contains $x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=0,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=0,w_{i}(e)=0$ for every other edge e on P.
  • Case 3: $\alpha_{q}=True$, $u_{p}$ contains $\neg x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=0,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=0,w_{i}(e)=0$ for every other edge e on P.
  • Case 4: $\alpha_{q}=False$, $u_{p}$ contains $\neg x_{q}$. In this case $w_{i}(e_{i_{2\cdot q-1}})=1,w_{i}(e_{i_{2\cdot n+2\cdot p-1}})=-1,w_{i}(e)=0$ for every other edge e on P.
In each of these cases, $z_{i}=0$. Therefore, for a path $P=(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$, where $e_{i_{1}}$ is adjacent to $A_{1}$ and $e_{i_{2\cdot(m+n)}}$ is adjacent to $Q_{m}$, we have $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=0$.
Conversely, let $P=(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$ be a path, such that $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=0$. The current goal is to choose values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ for $x_{1},x_{2},x_{3},...,x_{n}$, such that all clauses $u_{1},u_{2},...,u_{m}$ are satisfied. For $1\leq i\leq n$, let $\alpha_{i}=True$, if $e_{i_{2\cdot i-1}}=a_{i}$, and let $\alpha_{i}=False$, if $e_{i_{2\cdot i-1}}=d_{i}$. The next step is to check if all clauses from U are satisfied. Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause. If $e_{i_{2\cdot n+2\cdot j-1}}\in\{(P_{j},R_{j,1}),(P_{j},R_{j,2}),(P_{j},R_{j,3}),(P_{j},R_{j,4})\}$, then because $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=0$, either $u_{j}$ contains $x_{\alpha}$ and $e_{2\cdot\alpha-1}=a_{\alpha}$ , which means that $x_{\alpha}=True$ and $u_{j}$ is satisfied, or $u_{j}$ contains $\neg x_{\alpha}$ and $e_{2\cdot\alpha-1}=d_{\alpha}$, which means that $x_{\alpha}=False$ and $u_{j}$ is satisfied again. In both cases $u_{j}$ is satisfied after values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ are assigned. Cases when $e_{i_{2\cdot n+2\cdot j-1}}\in\{(P_{j},R_{j,5}),(P_{j},R_{j,6}),(P_{j},R_{j,7})\}$ can be resolved with similar reasoning, with $x_{\beta}$ or $x_{\gamma}$ being considered instead of $x_{\alpha}$. Therefore, after values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ were chosen, all clauses $u_{1},u_{2},...,u_{m}$ are satified.
Therefore, this process of obtaining an improper weight function problem from a 3-SAT problem is a polynomial transformation of a 3-SAT problem into an improper weight function problem, which makes an improper weight function problem an NP-complete one.
Proposition 3.
For an arbitrary graph $G(V,E)$ without loops and parallel edges, a sufficiently large prime number $p$ and an arbitrary natural $k$, $\chi_{MA(G)}(p^{k})$ is equal to the number of proper weight functions $W:E\rightarrow F_{p}^{k}$.
Proof.
Let $A$ be an arbitrary arrangement of hyperplanes in $R^{n}$ such that the equation of each hyperplane $A$ has integer coefficients. Let $A_{q}$ be a subset of all elements $F_{q}^{n}$ satisfying the equation of at least one of the hyperplanes $A$ modulo $p$, where $q=p^{k}$ is a prime power. The finite field method allows us to find the value of the characteristic polynomial of the arrangement $A$ at the point $q$ for sufficiently large $p$: $\chi_{A}(q)=|F_{q}^{n}\setminus A_{q}|$.
Let $q=p^{k}$ be a power of a prime number, and let $d$ be an irreducible polynomial of degree $k$ over $F_{p}$. Then there is a one-to-one correspondence between $F_{p}^{k}$ and $F_{q}$ that preserves addition and multiplication by elements of $F_{p}$: each element $(a_{1},a_{2},...,a_{k})$ from $F_{p}^{k}$ is associated with the element $a_{1}+a_{2}x+a_{3}x^{2}+....+a_{k}x^{k-1}$ in $F_{p}(x)/d\cong F_{q}$. This map induces a map from the set of functions $W:E\rightarrow F_{p}^{k}$ to the set of functions $\hat{W}:E\rightarrow F_{q}$. A proper function $W$ will be mappped to a function $\hat{W}$, which has the following property: $(\hat{W}(e_{1}),\hat{W}(e_{2}),..,\hat{W}(e_{n}))$ belongs to $F_{q}^{n}\setminus MA(G)_{q}$, due to (1) and the definition of $MA(G)$.
Thus, for a sufficiently large $p$ and an arbitrary $k$, $\chi_{MA(G)}(p^{k})$ is equal to the number of proper weight functions $W:E\rightarrow F_{p}^{k}$. ∎
An application of the improper weight function problem in cryptography
This chapter contains a description of a cryptosystem that uses an alternating weighted path problem, which is an extention of the improper weight function problem. Given a graph $G(V,E)$, a prime number $p$, a natural number $k$, a weight function $W:E\rightarrow F_{p}^{k}$ and a vector $(a_{1},...,a_{k})\in F_{p}^{k}$, the problem is to determine, if there is a sequence of edges $(e_{i_{1}},...,e_{i_{d}})$ that form a simple path or a simple even cycle in graph $G$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{k})$. Since we can use $(e_{i_{1}},...,e_{i_{d}})$ as a clue that would allow to solve this problem in polynomial time, the problem is NP. Moreover, we can reduce a general improper weight function problem to a specific alternating weighted path problem by picking $G,p,k,W$ as they were in the initial improper weight function problem and choosing $(a_{1},...,a_{k})={\bf 0}\in F_{p}^{k}$, making alternating weighted path problem an NP-hard problem. Since alternating weighted path problem is both NP and NP-hard, it’s NP-complete.
Despite the general alternating weighted path problem being NP-complete, there are special cases of this problem that can be solved fairly quickly. For example, let G be a complete graph $K_{m}$, let p be a large prime, $p>2\cdot 3^{m-1}$, and let $k=m$. Let $V=\{v_{1},...,v_{m}\}$ be a set of vertices of graph G. For an edge $e=(v_{q},v_{r}),q<r$, let the weight function $W(e)=(w_{1}(e),...,w_{m}(e))$ be defined the following way: $w_{t}(e)=0$, if $t\neq q,t\neq r$, $w_{t}(e)=3^{r-q-1}$, if $t=q$, and $w_{t}(e)=1$, if $t=r$. The key idea of the solution will be for a given $x\in F_{p}$ to either find an $1\leq i\leq m$ such that $x\equiv 3^{i-1}(mod\leavevmode\nobreak\ p)$ or $x\equiv p-3^{i-1}(mod\leavevmode\nobreak\ p)$, or find $1\leq i<j\leq m$, such that $x\equiv 3^{j}-3^{i}(mod\leavevmode\nobreak\ p)$ or $x\equiv p-3^{j}+3^{i}(mod\leavevmode\nobreak\ p)$. Note that because $p>2\cdot 3^{m-1}$, and because for any $1\leq i<j\leq m$ and $r\in\{0,1,3,...,3^{i-1}\},s\in\{0,1,3,...,3^{j-1}\}$ $3^{j}-s>3^{i}-r$, it is impossible to get multiple different answers for an $x$ as a result of this search. Also note, that since |$|\{3^{i-1}|1\leq i\leq m\}\cup\{p-3^{i-1}|1\leq i\leq m\}\cup\{3^{j}-3^{i}|1\leq i<j\leq m\}\cup\{p-3^{j}+3^{i}|1\leq i<j\leq m\}|=m(m+1)$, this search can be done in polynomial time.
For a vector $A=(a_{1},...,a_{m})\in F_{p}^{k}$, there is the following solution to this problem. The solution consists of m-1 steps, and the i-th step ends with either the conclusion, that the answer is “No”, or with the set of the edges $(v_{i},v_{j}),i<j$ that belong to a path or an even cycle $D=(e_{i_{1}},...,e_{i_{d}})$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$. In essense, this algoritm considers vertices of graph G one by one, from $v_{1}$ to $v_{m-1}$, tries to determine for each vertex, if it has adjacent edges that are part of $D$, and then adjusts the vector of weights, so that weights on edges of verices with lower indexes no longer affect the search of edges on D, adjacent to vertices with higher indexes.
Let’s describe the first step. If there is a path or an even cycle $D=(e_{i_{1}},...,e_{i_{d}})$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$, then there are 3 possible options. The first option is that $a_{1}=0$. In this case $D$ doesn’t contain any edges that are adjacent to $v_{1}$. The second option is that $a_{1}=3^{b},b<m$ or $a_{1}\equiv p-3^{b}$ (mod p). In this case $(v_{1},v_{b+2})$ is in $D$ and it is the only edge in $D$ that is adjacent to $v_{1}$. The third option is that $a_{1}\equiv 3^{b}-3^{c}$ (mod p), $b<m,c<m$. In this case both edges $(v_{1},v_{b+2})$ and $(v_{1},v_{c+2})$ are in $D$ and they are the only edges in $D$ that are adjacent to $v_{1}$. If $a_{1}$ doesn’t fit into any of these three options, then the answer to the alternating weighted path problem is “No”. Otherwise, the set of edges that are adjacent to $v_{1}$ and are in the path D, is stored, and the new vector of weights $B_{1}$ is obtained from $A$ in the following way. If $a_{1}=0$ then the stored set of edges is empty and $B_{1}=A$ If $a_{1}=3^{i}$ then the edge $(v_{1},v_{i})$ is stored and $B_{1}=A-W((v_{1},v_{i}))$. If $a_{1}\equiv p-3^{i}$ (mod p), then the edge $(v_{1},v_{i})$ is stored and $B_{1}=A+W((v_{1},v_{i}))$. If $a_{1}\equiv 3^{i}-3^{j}$ (mod p), then both edges $(v_{1},v_{i})$ and $(v_{1},v_{j})$ are stored and $B_{1}=A-W((v_{1},v_{i}))+W((v_{1},v_{j}))$. And this point the first step ends.
For the i-th step the process is the same, except $v_{i}$ is used instead of $v_{1}$, vector $B_{i-1}=(b_{1},...,b_{m})$, that we obtained at the end of $(i-1)$-th step, is used instead of A, $b_{i}$ is used instead of $a_{1}$, and a new vector $B_{i}$ is obtained as a result instead of $B_{1}$.
If all the steps are completed, and none of them ended with the conclusion that the answer is “No”, then the set of all stored edges is inspected. If all edges from this set form a simple path or a simple even cycle $D$, and it fits the condition $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$ , then the answer to the problem is “Yes”, otherwise the answer is “No”.
Now let’s describe a cryptosystem based on the alternating weighted path problem.
  • Public Key: Graph $G(V,E)$, prime number $p$, natural number $k$, and weight function $W:E\rightarrow F_{p}^{k}$.
  • Private Key: An invertible linear operator $M:F_{p}^{k}\rightarrow F_{p}^{k}$ that transforms weights on edges, such that an alternating weighted path problem transforms into an easy one.
  • Message Space: simple paths and simple even cycles of $G$.
  • Encryption: for a path or an even cycle $(e_{i_{1}},...,e_{i_{d}})$, compute $(a_{1},...,a_{k})=W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})$.
  • Decryption: apply operator $M$ to the received message $A=(a_{1},...,a_{k})$ and solve the easy alternating weighted path problem with weight function $W^{\prime}(e)=M(W(e))$ and the alternating sum of weights $B=M(A)$.
The example below shows a public key, a private key, and a weight function obtained after using private key.
Figure 7: Example
The important part of this cryptosystem is the superincreasing sequence in the easy version of the alternating weighted path problem. Similar approach is used in the Merkle-Hellman cryptosystem that is based on the knapsack problem [10]. In the Merkle-Hellman cryptosystem, the superincreasing sequence is disguised with a modular multiplication. This way of disguise is known to be vulnerable to attacks [11], [12]. In our cryptosystem, the superincreasing sequence is disguised with a linear operator, which is a broader method of disguise than the modular multiplication. Thus there might be a variation of this cryptosystem that is less vulnerable to attacks than knapsack-based cryptosystems.